OpenStudy (anonymous):

Simplify f(x) = (3x^2-4x+1)/(3x-1). Using complete sentences, explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f(1/3) is undefined. please show how u do it so i can understand

OpenStudy (johnweldon1993):

Can you factor the top?

OpenStudy (kenljw):

Factor numerator 3 x^2 - 4x + 1 = (-3 x + 1) (-1 x + 1) f(x) = (3 x -1) (x - 1)/(3 x -1) f(1) =0 f(0)=-1 f(1/3)=0/0 ? With Calculus L'Hospital rule f(1/3)=(6( (1/3) - 4)/3 = -2/3 this is a continuous function

OpenStudy (kenljw):

L'Hosptal Rule f(x) = N(x)/D(x) f(x1) = 0/0 or infinity/infinity f(x1) = N'(x1)/D'(x1) in the above problem 3x -1 do cancel out